GRAPHS AND SEPARABILITY PROPERTIES OF GROUPS

Rita Gitik

Abstract. A group G is LERF (locally extended residually finite) if for any finitely generated subgroup S of G and for any g ∉ S there exists a finite index subgroup S_{0} of G which contains S but not g. Using graph-theoretical methods we give algorithms for constructing finite index subgroups in amalgamated free products of groups with good separability properties. We prove that a free product of a free group and a LERF group amalgamated over a cyclic subgroup maximal in a free factor is LERF. The maximality condition cannot be removed, because adjunction of roots does not preserve property LERF. We also give short proofs of some old theorems about separability properties of groups, including a theorem of Brunner-Burns-Solitar that a free product of free groups amalgamated over a cyclic subgroup is LERF.

Typeset by AMS-TEX